# Elastic Constants and Bending Rigidities from Long-Wavelength Perturbation Expansions

### v1

@@ Appendix D Cluster expansion model for multipole and dielectric tensors
$\sum_{\alpha\in\mathbb{A}/\mathbb{G}_{S}}\sum_{\hat{s}\alpha\in\mathbb{G}_{S\alpha}}\mathfrak{R}_{IJ}(\hat{s})\mathcal{Q}_{\bm{\kappa}}^{J}(\alpha)=0.$ (140)
Unlike the DFPT or finite-difference scheme adopted in most ab initio codes where the rum rule of charge neutrality conditions is fulfilled as a post-processing step [60, 164], we explicitly consider it as an additional set of linear constraints when determining the number of irreducible parameters of multipole tensors. We expect this can further improve the accuracy of our fitted multipole tensors which will automatically satisfy the charge neutrality conditions.
@@
$\mathbb{B}(\bm{\kappa},\alpha I)=\frac{e}{\Omega}\frac{(-i)^{|\alpha|}}{\alpha!}\sum_{\hat{s}\alpha\in\mathbb{G}_{S\alpha}}\mathfrak{R}_{JI}(\hat{s})q_{J},$ (144)
where the composite row index $\bm{\kappa}$ is contained in the cluster $\alpha$ on the right-hand side, i.e. the first atom of each cluster. Also, the null space $\mathbb{N}$ has the shape of $N_{\bm{\mathcal{Q}}}\times N_{\bm{\mathcal{Q}}^{\prime}}$, where $N_{\bm{\mathcal{Q}}}$ and $N_{\bm{\mathcal{Q}}^{\prime}}$ are the number of total and irreducible parameters of multipole tensors, respectively. Then, an efficient least-square regression can be adopted to solve the linear problem in Eq. (143), and the solvability condition is that $\mathbb{B}^{\prime}$ needs to be full-rank, i.e. an overdetermined linear system. A simple estimation of the minimum number of configurations of phonon perturbation can be done via $N^{\mathrm{min}}_{\mathbf{q}}=N_{\bm{\mathcal{Q}}^{\prime}}/(6n_{\mathrm{a}})$, where the factor of 6 comes from the fact that each atom has three Cartesian components and the charge density response is a complex number. Even though we are using a set of random small $\mathbf{q}\to\mathbf{0}$ perturbations, it is still possible that some rows of $\mathbb{B}^{\prime}$ are linearly correlated, and hence we recommend to generate the number of $\mathbf{q}$ perturbations that is of 2 to 3 folds of $N^{\mathrm{min}}_{\mathbf{q}}$, in order to obtain the reliable and accurate solution of multipole tensors. Last but not least, the linear regression problem of dielectric screening function can be regarded as a special case of charge density response (i.e. for a single atom along one specific direction), and here we are not going to discuss it further.
Table 6: The dynamical Born effective charge $Z$ [$e$], quadrupole $Q$ [$e\cdot\mathrm{bohr}$], octupole $O$ [$e\cdot\mathrm{bohr}^{2}$], clamped-ion dielectric permittivity $\bm{\epsilon}^{\infty}$ and dielectric dispersion tensors $\bm{\epsilon}^{(4)}$ [bohr2] of BaTiO3, calculated by a least-square fit to the Taylor expansion of unscreened (or transverse) charge density response and dielectric screening function in Eqs. (43) and (II.4), respectively. We report without parentheses the results obtained with the PC prescription, while the corresponding LOC ones are provided in parentheses. Only symmetry-independent components are shown. Evidently, the relevant differences that cannot be attributed to numerical error arise only in the case of octupoles.
@@
$\epsilon^{(4)}_{xxxx}$27.855 (-27.859)
$\epsilon^{(4)}_{zzzz}$24.082 (-24.081)
$\epsilon^{(4)}_{yyyz}$2.142 (-2.142)
$\epsilon^{(4)}_{xxzz}$10.183 (-10.181)
To validate our approach for extracting multipole and dielectric tensors from a least-square linear regression of the cluster expansion model proposed here, we take the rhombohedral BaTiO3 as an example since it is the only material for which octupole interactions were reported in the literature [88]. For this comparison, we use their structure for BaTiO3 and keep every computational parameter the same as reported in Ref. [88]. A set of 20 $\mathbf{q}$ perturbations is randomly sampled over a small region around the zone center in arbitrary directions, with the magnitude of $\mathbf{q}$ ranging from 0.01 to 0.05 in reciprocal lattice unit. Then, the calculations of charge density response and dielectric function for those $\mathbf{q}$ points are then performed based on a recent implementation within the DFPT framework [166, 167], with the unscreened charge response obtained by switching off the macroscopic potential. Our results for the multipole and dielectric tensors of BaTiO3 are summarized in Table 5 and are in excellent agreement with Ref. [88]. Also, we note that the choice of the rhombohedral cell in Ref. [88] is not standard and gives rise to different symmetry-independent components, compared to the results shown in Table 6 (i.e. the one used in our elastic constant calculations). In general, when any linear regression technique is applied to get the expansion coefficient of a Taylor series, it is difficult to remove the cross talk among different orders, even if the system-dependent magnitude of correlation matrix is carefully chosen. There is a renormalization effect on the fitted quadrupole and octupole tensors from the higher-order terms that are not included in the expansion. Since a higher even order can only affect the lower even order (similarly for odd order), the obtained quadrupoles and octupoles should inherently contain the influence from hexadecapoles ($n=4$) and triacontadipoles ($n=5$), respectively. In order to obtain the unscreened and exact quadrupole and octupole tensors without those higher-order effects, we recommend to include those next-leading terms in the expansion, although this increases the number of DFPT calculations. Nevertheless, the exact values of mutipole tensors are not necessary for accurate phonon dispersion or short-circuit elastic constant calculations. In Table 5, we have performed the multipole expansion up to the hexadecapole term to ensure the obtained quadrupole tensor is unscreened and exact, which can be readily compared with those values from a direct linear response calculation in Ref. [88].
@@ Appendix E Influence of the choice of ionic charge representation in pseudopotentials
where $v(\mathbf{q})=4\pi e^{2}/|\mathbf{q}|^{2}$ and $V^{\textrm{ion,LOC}}_{\kappa\alpha}(\mathbf{q})$ is the local part of the pseudopotential. Clearly, the PC and LOC representations of ionic charges are inequivalent choices, having an impact on the determination of the multipolar expansion of charge density response. This problem has been discussed in App. B3 of Ref. [167], where Eqs. (145) and (146) are Eqs. (B29) and (B28), respectively. We report here, following the arguments of the cited reference, the proof that in semiconductors the choice of PC or LOC leads to different values for octupole tensors, while dipoles and quadrupoles should not be affected by the choice of the description of the ionic displacement.
@@
where $\epsilon^{-1}_{\textrm{L,PP}}(\mathbf{q})$ is the inverse dielectric function computed in presence of a pseudopotential, and the subscript “L” is used to distinguish it from the response function which relates the external to the full self-consistent potential. We indicate with $\overline{\rho}^{\textrm{PP,PC}}_{\kappa\alpha}(\mathbf{q})$ and $\overline{\rho}^{\textrm{PP,LOC}}_{\kappa\alpha}(\mathbf{q})$ the unscreened total charge density variations, where the bare atomic displacement is added to the self-consistently determined induced density via Eqs. (145) and (146), respectively, and the same notation is used for the screened responses. For simplicity, we consider the case of a cubic semiconductor which has the following relation holds:
@@
where we have defined $\rho^{\textrm{ion,PC}}_{\kappa\alpha}(\mathbf{q})=eV^{\textrm{ion,PC}}_{\kappa\alpha}(\mathbf{q})/v(\mathbf{q})$. As discussed in App. C of Ref. [157], one can demonstrate
@@ Appendix F Computational details
First-principles calculations are performed using the Quantum ESPRESSO package [104, 105], with the norm-conserving pseudopotentials from PseudoDojo [170]. We employ the Perdew, Burke and Ernzerhof’s (PBE) [171] parametrization of the generalized gradient approximation (GGA) for silicon, NaCl and graphene, and the Perdew-Wang’s parametrization of the local density approximation (LDA) [172] for BaTiO3, to describe the exchange-correlation effect of electrons. In the case of GaAs, the revision of PBE functional for solid (PBEsol) [173] is adopted to overcome the infamous problem of metallic-like screening with the standard PBE and have better lattice parameters compared to experimental values. For silicon, NaCl and GaAs, the electronic $\mathbf{k}$ grid is chosen to be 12$\times$12$\times$12, while a smaller 10$\times$10$\times$10 grid is used for BaTiO3, and their plane-wave energy cutoffs are 40, 160, 90 and 100 Ry, respectively. To simulate graphene as a 2D system, the Coulomb cutoff technique [122] is used to remove the long-range electrostatic interactions along the vacuum direction, using the Marzari-Vanderbilt smearing [174] with a smearing of 0.02 Ry, where the electronic $\mathbf{k}$ grid and plane-wave energy cutoff are set to 48$\times$48$\times$1 and 100 Ry, respectively. For other 2D materials including h-BN, MoS2 and InSe, we use the same primitive cell structure and computational parameters as described in Ref. [98]. All of structural optimizations are carried out based on the convergence thresholds as the pressure and forces smaller than $10^{-3}$ kbar and $10^{-4}$ Ry/bohr. Then, the DFPT simulations are performed to calculate the dynamical matrices on a set of phonon grids, from which the IFCs used in this work are obtained by taking the inverse Fourier transform. We modified the matdyn code of the Quantum ESPRESSO software [104, 105] and implemented Eq. (40) to obtain the elastic constants and Eq. (78) for the bending rigidities, where the macroscopic long-range electrostatics has been removed. For the bulk materials considered in this study, we perform a least-square fit to the Taylor expansion of the unscreened (or transverse) charge density response and dielectric screening function to a phonon perturbation, to extract the multipole and dielectric tensors up to the third and fourth orders as in Eqs. (43) and (II.4), respectively (see App. D for details). We generate a list of $\mathbf{q}$ points (5 to 10 depending on crystal symmetry) along random directions with the magnitude ranging from 0.01 to 0.05 of the reciprocal lattice constant, and the unscreened charge density response after removing the macroscopic potential as well as the dielectric function for those $\mathbf{q}$ points are calculated within the DFPT framework based on recent developments [167, 118, 175, 166]. The electronic grid for GaAs is further increased to 30$\times$30$\times$30 to accurately predict the dielectric properties of the system. The dynamical quadrupole tensors of 2D h-BN, MoS2 and InSe are directly taken from Ref. [98] which were computed using the abinit software [158, 159].
@@
For the finite-difference calculations of elastic constants of bulk solids, we use the strain-stress relation in Eq. (1) as implemented in the thermo_pw [55] code which externally relies on Quantum ESPRESSO subroutines. A few independent strains with a magnitude of 0.005 are applied to deform the system through the specific directions, according to its Laue class and crystal symmetry. Then, several self-consistent field calculations are performed to obtain the corresponding stress tensor of those deformations, with the allowed relaxation of ionic positions to the strain perturbation. The elastic stiffness tensor as the first-order derivative of stress with respect to strain is calculated numerically by finite difference.

### v2

@@ Appendix D Cluster expansion model for multipole and dielectric tensors
$\sum_{\alpha\in\mathbb{A}/\mathbb{G}_{S}}\sum_{\hat{s}\alpha\in\mathbb{G}_{S\alpha}}\mathfrak{R}_{IJ}(\hat{s})\mathcal{Q}_{\bm{\kappa}}^{J}(\alpha)=0,$ (140)
where $\mathbb{A}$ stands for a set of all possible clusters, and $\mathbb{A}/\mathbb{G}_{S}$ is the orbit space that the representative clusters live in. Unlike the DFPT or finite-difference scheme adopted in most ab initio codes where the sum rule of charge neutrality conditions is fulfilled as a post-processing step [64, 173], we explicitly consider it as an additional set of linear constraints when determining the number of irreducible parameters of multipole tensors. We expect this can further improve the accuracy of our fitted multipole tensors which will automatically satisfy the charge neutrality conditions.
@@
$\mathbb{B}(\bm{\kappa}^{\prime},\alpha I)=\frac{e}{\Omega}\frac{(-i)^{|\alpha|}}{\alpha!}\sum_{\hat{s}\alpha\in\mathbb{G}_{S\alpha}}\mathfrak{R}_{JI}(\hat{s})q_{J},$ (144)
where the composite row index $\bm{\kappa}^{\prime}$ is contained in the cluster $\hat{s}\alpha$ on the right-hand side, i.e. the first atom of each cluster. Also, the null space $\mathbb{N}$ has the shape of $N_{\bm{\mathcal{Q}}}\times N_{\bm{\mathcal{Q}}^{\prime}}$, where $N_{\bm{\mathcal{Q}}}$ and $N_{\bm{\mathcal{Q}}^{\prime}}$ are the number of total and irreducible parameters of multipole tensors, respectively. Then, an efficient least-squares regression can be adopted to solve the linear problem in Eq. (143), and the solvability condition is that $\mathbb{B}^{\prime}$ needs to be full-rank, i.e. an overdetermined linear system. A simple estimation of the minimum number of configurations of phonon perturbation can be done via $N^{\mathrm{min}}_{\mathbf{q}}=N_{\bm{\mathcal{Q}}^{\prime}}/(6n_{\mathrm{a}})$, where the factor of 6 comes from the fact that each atom has three Cartesian components and the charge density response is a complex number. Even though we are using a set of random small $\mathbf{q}\to\mathbf{0}$ perturbations, it is still possible that some rows of $\mathbb{B}^{\prime}$ are linearly correlated, and hence we recommend to generate the number of $\mathbf{q}$ perturbations that is of 2 to 3 folds of $N^{\mathrm{min}}_{\mathbf{q}}$, in order to obtain the reliable and accurate solution of multipole tensors. Last but not least, the linear regression problem of dielectric screening function can be regarded as a special case of charge density response (i.e. for a single atom along one specific direction), and here we are not going to discuss it further.
Table 6: The dynamical Born effective charge $Z$ [$e$], quadrupole $Q$ $e~\mathrm{bohr}$], octupole $O$ $e~\mathrm{bohr}^{2}$], clamped-ion dielectric permittivity $\bm{\epsilon}^{\infty}$ and dielectric dispersion tensors $\bm{\epsilon}^{(4)}$ [bohr2] of BaTiO3, calculated by a least-squares fit to the Taylor expansion of unscreened (or transverse) charge density response and dielectric screening function in Eqs. (43) and (II.4), respectively. We report without parentheses the results obtained with the PC prescription, while the corresponding LOC ones are provided in parentheses. Only symmetry-independent components are shown. Evidently, the relevant differences that cannot be attributed to numerical error arise only in the case of octupoles.
@@
$\epsilon^{(4)}_{xxxx} -$27.855 (-27.859)
$\epsilon^{(4)}_{zzzz} -$24.082 (-24.081)
$\epsilon^{(4)}_{yyyz} -$2.142 (-2.142)
$\epsilon^{(4)}_{xxzz} -$10.183 (-10.181)
To validate our approach for extracting multipole and dielectric tensors from a least-squares linear regression of the cluster expansion model proposed here, we take the rhombohedral BaTiO3 as an example since it is the only material for which octupole interactions were reported in the literature [95]. For this comparison, we use their structure for BaTiO3 and keep every computational parameter the same as reported in Ref. [95]. A set of 20 $\mathbf{q}$ perturbations is randomly sampled over a small region around the zone center in arbitrary directions, with the magnitude of $\mathbf{q}$ ranging from 0.01 to 0.05 in reciprocal lattice unit. Then, the calculations of charge density response and dielectric function for those $\mathbf{q}$ points are then performed based on a recent implementation within the DFPT framework [175, 176], with the unscreened charge response obtained by switching off the macroscopic potential. Our results for the multipole and dielectric tensors of BaTiO3 are summarized in Table 5 and are in excellent agreement with Ref. [95]. Also, we note that the choice of the rhombohedral cell in Ref. [95] is not standard and gives rise to different symmetry-independent components, compared to the results shown in Table 6 (i.e. the one used in our elastic constant calculations). In general, when any linear regression technique is applied to get the expansion coefficient of a Taylor series, it is difficult to remove the crosstalk among different orders, even if the system-dependent magnitude of correlation matrix is carefully chosen. There is a renormalization effect on the fitted quadrupole and octupole tensors from the higher-order terms that are not included in the expansion. Since a higher even order can only affect the lower even order (similarly for odd order), the obtained quadrupoles and octupoles should inherently contain the influence from hexadecapoles ($n=4$) and triacontadipoles ($n=5$), respectively. In order to obtain the unscreened and exact quadrupole and octupole tensors without those higher-order effects, we recommend to include those next-leading terms in the expansion, although this increases the number of DFPT calculations. Nevertheless, the exact values of multipole tensors are not necessary for accurate phonon dispersion or short-circuit elastic constant calculations. In Table 5, we have performed the multipole expansion up to the hexadecapole term to ensure the obtained quadrupole tensor is unscreened and exact, which can be readily compared with those values from a direct linear response calculation in Ref. [95].
Levels $C_{11}$ $C_{12}$ $C_{13}$ $C_{14}$ $C_{33}$ $C_{44}$
PC  LOC PC  LOC PC  LOC PC  LOC PC  LOC PC  LOC
$D$ 242.90 242.90 81.37 81.37 48.29 48.29 70.87 70.87 293.84 293.84 25.43 25.43
DD+DQ+$Q$ 294.68 294.67 86.31 86.31 60.79 60.80 44.72 44.72 299.96 299.98 56.67 56.69
[…]
@@ Appendix E Influence of the choice of ionic charge representation in pseudopotentials
where $v(\mathbf{q})=4\pi e^{2}/|\mathbf{q}|^{2}$ and $V^{\textrm{ion,LOC}}_{\kappa\alpha}(\mathbf{q})$ is the local part of the pseudopotential. Clearly, the PC and LOC representations of ionic charges are inequivalent choices, having an impact on the determination of the multipolar expansion of charge density response. This problem has been discussed in Appendix B3 of Ref. [176], where Eqs. (145) and (146) are Eqs. (B29) and (B28), respectively. We report here, following the arguments of the cited reference, the proof that in semiconductors the choice of PC or LOC leads to different values for octupole tensors, while dipoles and quadrupoles should not be affected by the choice of the description of the ionic displacement.
@@
where $\epsilon^{-1}_{\textrm{L,PP}}(\mathbf{q})$ is the inverse dielectric function computed in presence of a pseudopotential, and the subscript “L” is used to distinguish it from the response function which relates the external to the full self-consistent potential. We indicate with $\overline{\rho}^{\textrm{PP,PC}}_{\kappa\alpha}(\mathbf{q})$ and $\overline{\rho}^{\textrm{PP,LOC}}_{\kappa\alpha}(\mathbf{q})$ the unscreened total charge density variations, where the bare atomic displacement is added to the self-consistently determined induced density via Eqs. (145) and (146), respectively, and the same notation is used for the screened responses. For simplicity, we consider the case of a cubic semiconductor, where the following relation holds:
@@
where we have defined $\rho^{\textrm{ion,PC}}_{\kappa\alpha}(\mathbf{q})=eV^{\textrm{ion,PC}}_{\kappa\alpha}(\mathbf{q})/v(\mathbf{q})$. As discussed in Appendix C of Ref. [165], one can demonstrate
@@ Appendix F Computational details
First-principles calculations are performed using the Quantum ESPRESSO package [111, 112], with the norm-conserving pseudopotentials from PseudoDojo [179]. We employ the Perdew, Burke, and Ernzerhof’s (PBE) [180] parametrization of the generalized gradient approximation (GGA) for silicon, NaCl and graphene, and Perdew-Wang’s parametrization of the local density approximation (LDA) [181] for BaTiO3, to describe the exchange-correlation effect of electrons. In the case of GaAs, the revision of PBE functional for solid (PBEsol) [182] is adopted to overcome the infamous problem of metallic-like screening with the standard PBE and have better lattice parameters compared to experimental values. For silicon, NaCl, and GaAs, the electronic $\mathbf{k}$ grid is chosen to be $\times12$$\times$12, while a smaller $\times10$$\times$10 grid is used for BaTiO3, and their plane-wave energy cutoffs are 40, 160, 90, and 100 Ry, respectively. To simulate graphene as a 2D system, the Coulomb cutoff technique [133] is used to remove the long-range electrostatic interactions along the vacuum direction, using the Marzari-Vanderbilt smearing [183] with a smearing of 0.02 Ry, where the electronic $\mathbf{k}$ grid and plane-wave energy cutoff are set to $\times48$$\times$1 and 100 Ry, respectively. For other 2D materials including h-BN, MoS2 and InSe, we use the same primitive cell structure and computational parameters as described in Ref. [105]. All of structural optimizations are carried out based on the convergence thresholds as the pressure and forces smaller than $10^{-3}$ kbar and $10^{-4}$ Ry/bohr. Then, the DFPT simulations are performed to calculate the dynamical matrices on a set of phonon grids, from which the IFCs used in this work are obtained by taking the inverse Fourier transform. We modified the matdyn code of the Quantum ESPRESSO software [111, 112] and implemented Eq. (40) to obtain the elastic constants and Eq. (78) for the bending rigidities, where the macroscopic long-range electrostatics has been removed. For the bulk materials considered in this study, we perform a least-squares fit to the Taylor expansion of the unscreened (or transverse) charge density response and dielectric screening function to a phonon perturbation, to extract the multipole and dielectric tensors up to the third and fourth orders as in Eqs. (43) and (II.4), respectively (see Appendix D for details). We generate a list of $\mathbf{q}$ points (five to ten depending on crystal symmetry) along random directions with the magnitude ranging from 0.01 to 0.05 of the reciprocal lattice constant, and the unscreened charge density response after removing the macroscopic potential as well as the dielectric function for those $\mathbf{q}$ points are calculated within the DFPT framework based on recent developments [176, 129, 184, 175]. The electronic grid for GaAs is further increased to $\times30$$\times$30 to accurately predict the dielectric properties of the system. The dynamical quadrupole tensors of 2D h-BN, MoS2, and InSe are directly taken from Ref. [105] which were computed using the abinit software [166, 167].
@@
For the finite-difference calculations of elastic constants of bulk solids, we use the strain-stress relation in Eq. (1) as implemented in the thermo_pw [59] code, which externally relies on Quantum ESPRESSO subroutines. A few independent strains with a magnitude of 0.005 are applied to deform the system through the specific directions, according to its Laue class and crystal symmetry. Then, several self-consistent field calculations are performed to obtain the corresponding stress tensor of those deformations, with the allowed relaxation of ionic positions to the strain perturbation. The elastic stiffness tensor as the first-order derivative of stress with respect to strain is calculated numerically by finite difference.